OpenStudy (anonymous):

Let f(t)=3(2)^t. Find the value m such that f(m)=500.

OpenStudy (anonymous):

Please Help! I'm incredibly lost right now.

OpenStudy (campbell_st):

you need to find the value of t that will give 500 so equate the 2 equations 500 = 3(2)^t 500/3 = 2^t take the logs of both sides , log 10 or log e I'll use base e ln(500/3) = ln(2^t) using log laws... the right hand side can be rewritten ln(500/3) = t x ln(2) so t = ln(500/3)/ln(2) evaluating and you should get the correct answer